### v2

# Coprime Bivariate Bicycle Codes and their Properties

Abstract
This work (1) proposes a novel numerical algorithm to accelerate the search process for good Bivariate Bicycle (BB) codes and (2) defines a new subclass of BB codes suitable for quantum error correction. The proposed acceleration search algorithm reduces the search space by excluding some equivalent codes from the search space, as well as setting thresholds to drop bad codes at an early stage. A number of new BB codes found by this algorithm are reported. The proposed subclass of BB codes employs coprimes to construct groups via polynomials as the basis for the BB code, rather than using the standard BB codes with unconstrained constructors. In contrast to vanilla BB codes, where parameters remain unknown prior to code discovery, the rate of the proposed code can be determined beforehand by specifying a factor polynomial as an input to the numerical search algorithm. Using this coprime BB construction, we found a number of surprisingly short to medium-length codes that were previously unknown.
1 Introduction
Quantum information is susceptible to errors during storage and operation. As the number of qubits in a quantum circuit increases, so does the frequency of errors. Therefore, quantum Error Correction (QEC) is the cornerstone of advancing from the current noisy intermediate-scale quantum (NISQ) era to the next era of fault-tolerant quantum (FTQC) computing. Typically, a QEC code is characterized by a 3-tuple $[[n,k,d]]$, signifying that the code utilizes $n$ physical qubits to encode $k$ logical qubits and can correct up to $\lfloor d/2\rfloor$ errors. This encoding process introduces additional overhead, which is measured by the code rate $k/n$. Among the various QEC codes, quantum Low-Density Parity-Check (qLDPC) codes stand out [1, 2, 3, 4] due to their lower-weight stabilizers, low overhead, and high thresholds. As a special case of qLDPC codes, the surface codes, which also feature low-weight stabilizers, are the most commonly used codes in quantum computing. However, surface codes require significant overhead. For example, the rotated surface code has parameters $[[L^{2},1,L]]$, meaning that it requires $L^{2}$ physical qubits to protect one logical qubit.
In contrast, research has demonstrated the existence of “good” qLDPC codes [3] indicating that qLDPC codes can have $k$ and $d$ scaling linearly as $n$ grows. However, having asymptotically good LDPC codes does not necessarily mean having better parameters than codes designed for short to medium lengths. Moreover, the structure of codes can greatly influence the design of hardware and vice versa. Therefore, for practical purposes, people started to seek qLDPC codes that are easy to implement on hardware and have good finite-length performance. In particular, one such qLDPC code, the BB code [4], has received much attention because they have high thresholds, toric layout, and can be embedded on two planes. Benefiting from the quasi-cyclic and two-thickness properties of BB codes, recent papers have established the feasibility of BB codes on different architectures, including cold atoms [5, 6] and superconducting [7]. In particular, Viszlai et al. [5] showed that the quasi-cyclic structure of BB codes can be easily implemented using atom array acousto-optic deflectors (AOD) movements. However, constructing BB codes is a time-consuming process because one has to search for combinations of polynomials to construct a code. Moreover, the parameters of codes can not be guaranteed.
This challenge motivated us to develop an algorithm that accelerates the search for good BB codes and even constructs codes with pre-determined parameters. In this work, we propose a fast numerical algorithm to search for good BB codes by excluding certain polynomial combinations. In addition, a new construction of BB codes is proposed that allows us to customize the code rate before performing a search, much in contrast to prior search techniques that identified the rate only after returning a new code as a search result. This new method involves selecting two coprime numbers and a factor polynomial, leading us to name this subclass coprime-BB codes.
The remainder of this paper is organized as follows: Section 2 covers the preliminaries of BB codes. Section 3 details the proposed fast search algorithm. Section 4 explains the construction of coprime-BB codes. Section 6 presents the simulation results and compares our proposed codes. Lastly, Section 7 introduces related work and discusses potential future improvements.
2 Preliminaries
2.1 Stabilizer Codes and CSS Codes
Let $P_{n}=\{I,X,Y,Z\}^{\otimes n}$ be the set of Pauli strings acting on $n$ qubits. Any Pauli string can be decomposed into the product of two Pauli strings by disregarding the global phase. One of them only contains $I$ and $X$, while the other only contains $I$ and $Z$. Both of these strings can be represented by length $n$ binary strings:
$P=X^{\bm{x}}Z^{\bm{z}}=\bigotimes_{i=1}^{n}X^{x_{i}}\bigotimes_{i=1}^{n}Z^{z_{i}}.$ (1)
Thus, we can use a length $2n$ binary string $(\bm{x}|\bm{z})$ to express a Pauli string $P$ in a binary symplectic representation. Since $X$ and $Z$ anti-commute, the commutativity for two Pauli strings $P_{1},P_{2}$ is equivalent to the orthogonality of their binary symplectic representations $(\bm{x}_{1}|\bm{z}_{1}),(\bm{x}_{2}|\bm{z}_{2})$, as shown below:
$[P_{1},P_{2}]=0\iff\bm{x}_{1}\cdot\bm{z}_{2}+\bm{z}_{1}\cdot\bm{x}_{2}=0$ (2)
Let $S$ be an abelian group whose elements are in $P_{n}$ and such that $-I^{\otimes n}\notin S$. The code defined by $S$ has a code space consisting of the common $+1$-eigenspace of the elements in $S$. This group $S$ is known as the stabilizer group, and its elements are called stabilizers. An $[[n,k,d]]$ stabilizer code can be defined by $n-k$ independent stabilizers, which can generate the entire stabilizer group $S$. The code can also be represented by an $(n-k)\times 2n$ binary parity-check matrix, where each row of the matrix corresponds to a stabilizer generator in binary symplectic form. The minimum distance $d$ is the minimum weight (the number of non-identity operations) of a Pauli string $E$ such that $E\notin S$ but commutes with all stabilizers in $S$. An error with a weight greater than $d$ could cause a logical error.
Calderbank-Shor-Steane(CSS) codes, as an important class of stabilizer codes, have the property that each generator of the stabilizer group is either of the $X$-type or $Z$-type. This characteristic allows us to correct errors by correcting $Z$ errors and $X$ errors separately. The parity-check matrix of CSS codes has the form of
$H_{CSS}=\begin{pmatrix}H_{X}&\bm{0}\\ \bm{0}&H_{Z}\end{pmatrix}\leavevmode\nobreak\ ,$ (3)
where $H_{X}$ and $H_{Z}$ are binary matrices that correspond to the $X$-type and $Z$-type stabilizer generators, respectively. Since all stabilizers commute with each other, it follows directly that for a CSS code, $H_{X}H_{Z}^{T}=0$, as indicated by Eq. (2).
2.2 Bivariate Bicycle Codes
Let $S_{m}$ be a circulant permutation matrix of size $m$, defined as
$S_{m}=I_{m}>>1,$ (4)
where ">>much-greater-than$>$ >" denotes the right cyclic shift. For example,
$S_{3}=\begin{bmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{bmatrix}.$ (5)
By defining $x=S_{l}\otimes I_{m}$ and $y=I_{l}\otimes S_{m}$, it is easy to verify that $xy=yx$ using the mixed-product property of the Kronecker product.
Bivariate bicycle codes can be defined by two polynomials, $a(x,y)$ and $b(x,y)$, with corresponding matrices $A$ and $B$, respectively. This construction allows us to treat the product of these matrices as the product of polynomials. Consequently, the matrices $A,B$ commute since their product can be written as the product of polynomials $x,y$, and the multiplication of polynomials is commutative. Additionally, from the definition of $x$ and $y$, we observe that $x^{l}=y^{m}=I_{lm}$. Thus, the polynomials $a(x,y)$ and $b(x,y)$ belong to the polynomial ring $\mathbb{F}_{2}[x,y]/(x^{l}+1,y^{m}+1)$.
Given integers $l,m$, polynomials $a(x,y),b(x,y)$ and their matrix representation $A,B$, respectively, the parity check matrices for BB code are defined as
$\begin{split}H_{X}&=[A|B]\\ H_{Z}&=[B^{T}|A^{T}]\end{split}$ (6)
and it meets the CSS condition as $H_{X}H_{Z}^{T}=AB+BA=2AB=\bm{0}$ in $GF(2)$.
In [4], the authors restricted the polynomials to shapes of
$\begin{split}a(x,y)&=x^{a}+y^{b}+y^{c}\\ b(x,y)&=y^{d}+x^{e}+x^{f}.\end{split}$ (7)
Each polynomial has three terms, resulting in a weight of six per stabilizer. Therefore, in matrix form, we can write $A=A_{1}+A_{2}+A_{3}$ and $B=B_{1}+B_{2}+B_{3}$. Besides, we have $A^{T}=A_{1}^{T}+A_{2}^{T}+A_{3}^{T}=A_{1}^{-1}+A_{2}^{-1}+A_{3}^{-1}$ as $A_{i}$ is the power of $x$ or $y$, which are permutation matrices. For arbitrary permutation matrix $P$, it holds that $P^{T}=P^{-1}$. Similarly, we have $B^{T}=B_{1}^{-1}+B_{2}^{-1}+B_{3}^{-1}$. It is well known [8] that for any CSS code, $k=n-\text{rank}(H_{X})-\text{rank}(H_{Z})$. But for BB codes, according to Lemma 1 in [4], $\text{rank}(H_{X})=\text{rank}(H_{Z})$, so this expression simplifies to
$k=2lm-2\text{rank}(H_{X}).$ (8)
3 Numerical Acceleration of Searches
3.1 Code equivalence
In [4], BB codes are obtained through a numerical search. To accelerate this search, we reduce the search space by eliminating some codes with the same $n,k,d$ parameters. It is straightforward to prove that the following four codes
$\begin{split}C_{1}:H_{X}&=[A|B],H_{Z}=[B^{T}|A^{T}]\\ C_{2}:H_{X}&=[A^{T}|B^{T}],H_{Z}=[B|A]\\ C_{3}:H_{X}&=[B|A],H_{Z}=[A^{T}|B^{T}]\\ C_{4}:H_{X}&=[B^{T}|A^{T}],H_{Z}=[A|B]\end{split}$ (9)
share the same parameters, allowing us to search within only one class of these codes.
Proof: According to Lemma 1 in [4], every bivariate bicycle code has the same distance over $X$ or $Z$ and $\text{rank}(H_{X})=\text{rank}(H_{Z})$. Let us assume the distances and dimensions of $C_{1},\dots,C_{4}$ are $d_{1},\dots,d_{4}$ and $k_{1},\dots,k_{4}$. For CSS codes, $k=n-\text{rank}(H_{X})-\text{rank}(H_{Z})$ and it is easy to see that $k_{1}=k_{4}$ and $k_{2}=k_{3}$ since one can get one code from another by swapping $H_{X}$ and $H_{Z}$. Thus, it is sufficient to prove that
  • $d_{1}=d_{4}$ and $d_{2}=d_{3}$.
  • $k_{1}=k_{2}$ and $d_{1}=d_{2}$.
to ensure $C_{1},\dots,C_{4}$ has the same parameters.
Let $\bm{f}$ be the binary string of an arbitrary logical $X$ operator of $C_{1}$. By the definition of logical operators, we have
$[B^{T}|A^{T}]\bm{f}^{T}=\bm{0},$ (10)
indicating that $Z^{\bm{f}}$ is a logical $Z$ operator of $C_{4}$. Hence $d_{1}\geq d_{4}$. Similarly, we have $d_{4}\geq d_{1}$. Therefore, $d_{4}=d_{1}$. Using the same reasoning, we can prove that $C_{2}$ and $C_{3}$ have the same distance.
Next, we will prove that $C_{1}$ and $C_{2}$ have the same distance. Let $C_{l}$ be the anti-diagonal matrix of size $l\times l$. We have $C_{l}C_{l}=I$ and $C_{l}MC_{l}=M^{T}$ for any given matrix $M$ of the same size. As per
$[A^{T}|B^{T}]=C_{lm}[A|B]{\begin{bmatrix}C_{lm}&0\\ 0&C_{lm}\end{bmatrix},}$ (11)
we know $[A^{T}|B^{T}]$ and $[A|B]$ have the same rank and $k_{1}=k_{2}$. Let $Z^{\bm{p}}$ be an arbitrary logical $Z$ operator of $C_{2}$, where $\bm{p}=(\bm{\alpha}|\bm{\beta})$ is a length $n$ binary vector and $\bm{\alpha},\bm{\beta}$ are binary vectors of length $n/2$. By definition, we have $[A^{T}|B^{T}]\bm{p}^{T}=A\bm{\alpha}^{T}+B\bm{\beta}^{T}=\bm{0}$. I.e.,
$C_{lm}AC_{lm}\bm{\alpha}^{T}+C_{lm}BC_{lm}\bm{\beta}^{T}=\bm{0}.$ (12)
Recall that $C_{lm}C_{lm}=I$. By multiplying both sides with $C_{lm}$, we get
$AC_{lm}\bm{\alpha}^{T}+BC_{lm}\bm{\beta}^{T}=\bm{0},$ (13)
and we know that $(\bm{\alpha}C_{lm}^{T}|\bm{\beta}C_{lm}^{T})$ is a logical $Z$ operator of $C_{1}$ with the same weight of $\bm{p}$. Therefore, $d_{1}\leq d_{2}$ and, similarly, $d_{2}\leq d_{1}$. Thus, $d_{2}=d_{1}$.∎
We note that these two codes $C_{1}:H_{X}=[A|B],H_{Z}=[B^{T}|A^{T}]$ and $C_{5}:H_{X}=[A^{T}|B],H_{Z}=[B^{T}|A]$ do not always have the same parameters. For example, when $l=6,m=12$, the code constructed by $a(x,y)=x^{4}+y^{2}+y^{6}$ and $b(x,y)=y^{5}+x^{3}+x^{4}$ is a $[[144,8,10]]$ code whereas the code constructed by $a(x,y)=x^{2}+y^{6}+y^{10}$ and $b(x,y)=y^{5}+x^{3}+x^{4}$ is a $[[144,8,8]]$ code.
3.2 Code Selection
After enumerating the polynomials, we construct the parity-check matrices and evaluate the parameters $k$ and $d$ of the code. The dimension $k$ can be easily calculated using Eq. (8). However, calculating the distance $d$ is computationally intensive using mixed integer programming (MIP). Instead, we estimate the upper bound of distance $\hat{d}$ during the search using decoding algorithms; if $\hat{d}$ is low, we can discard the code. Besides, if a code has a disconnected Tanner graph, the distance of this code will not be high. Therefore, we will exclude these codes before evaluating their distance upper bounds. After completing the search, we can select codes with high $k$ and $\hat{d}$ to further validate their exact distances using MIP.
As mentioned above, BB codes provide the same distance for $X$ and $Z$ errors. Therefore, it is sufficient to find the $X$ distance to determine the distance of the code. Given a random logical $Z$ operator, there exists a logical $X$ operator that anti-commutes with it and commutes with all $Z$ stabilizers. Let $\bm{u}$ be the binary vector form of a random logical $Z$ operator. Then we have
$d_{X}\leq\min\text{wt}(\bm{v})\text{ s.t. }\bm{u}\bm{v}^{T}=1,H_{Z}\bm{v}^{T}=\bm{0}.$ (14)
$d_{X}\leq\min\text{wt}(\bm{v})\text{ s.t. }\begin{bmatrix}H_{Z}\\ \bm{u}\end{bmatrix}\bm{v}^{T}=\begin{bmatrix}\bm{0}\\ 1\end{bmatrix}$ (15)
This condition is equivalent to decoding a code with an enlarged parity-check matrix $H^{\prime}=\begin{bmatrix}H_{Z}\\ \bm{u}\end{bmatrix}$ and syndrome $\bm{s}=\begin{bmatrix}\bm{0}\\ 1\end{bmatrix}$. In [4], the authors used the BP-OSD [9, 10] algorithm to perform the decoding. The distance can be tightly bounded after multiple rounds of iterations. Since we are finding codes with high distances, this process can be aborted if the upper bound is lower than the desired distance $\tau_{d}$.
The search algorithm is given in Algorithm 1, and some of the codes found using this algorithm are listed in Table 1. It is important to note that, although our search focused on codes with weight-6 stabilizers of the form specified in Eq.(7), the algorithm can be adapted to search for codes with different weights or forms by modifying its input parameters.
Data: $l,m,\tau_{k},\tau_{d}$
Result: codes of parameters $[[2lm,k,d]]$
Generate all polynomial pairs of the specified form $L\leftarrow[(a_{1}(x,y),b_{1}(x,y)),...]$;
Remove codes with the same parameters: $L^{\prime}\leftarrow\textnormal{{remove\_equivalent}}(L)$;
for $i\leftarrow 1$ to $|L^{\prime}|$ do
[…]
             $k\leftarrow 2lm-2\textnormal{{rank}}(H_{X})$;
             if $k<\tau_{k}$ then
                   continue ;
             else
                   $d\leftarrow\textnormal{{distance\_upperbound}}(H_{X},H_{Z},\tau_{d})$;
             end if
       else
             continue ;
       end if
end for
Algorithm 1 An algorithm to search for BB codes
Table 1: Novel codes found by Algorithm 1
$l$ $m$ $a(x,y)$ $b(x,y)$ $[[n,k,d]]$
3 9 $1+y^{2}+y^{4}$ $y^{3}+x^{1}+x^{2}$ $[[54,8,6]]$
7 7 $x^{3}+y^{5}+y^{6}$ $y^{2}+x^{3}+x^{5}$ $[[98,6,12]]$
3 21 $1+y^{2}+y^{10}$ $y^{3}+x+x^{2}$ $[[126,8,10]]$
5 15 $1+y^{6}+y^{8}$ $y^{5}+x+x^{4}$ $[[150,16,8]]$
3 27 $1+y^{10}+y^{14}$ $y^{12}+x+x^{2}$ $[[162,8,14]]$
6 15 $x^{3}+y+y^{2}$ $y^{6}+x^{4}+x^{5}$ $[[180,8,16]]$
4 Coprime Bivariate Bicycle Codes
Based on the commutativity of matrices $x$ and $y$, one can choose different polynomial forms and construct valid CSS codes. But we need to perform an extensive search to find codes with good $k,d$ using the polynomial of form in Eq. (7). Here, we propose a different construction that can give us codes with pre-determined $k$.
Following the definition of BB codes, let $l,m$ be two coprime numbers and $\pi=xy$. The BB code defined by polynomials $a(x,y)$ and $b(x,y)$ can be expressed by univariate polynomials $a(\pi)$ and $b(\pi)$. Let $g(\pi)=\text{GCD}(a(\pi),b(\pi),\pi^{lm}+1)$, the BB code defined by $a(\pi)$ and $b(\pi)$ will have dimension
$k=2\deg g(\pi).$ (16)
Proof: We note that $\langle x\rangle$ and $\langle y\rangle$ are cyclic groups of order $l$ and $m$, respectively. It is easy to verify that $\langle xy\rangle$ is also a cyclic group of order $lm$, and any monomial in $\{x^{i}y^{j}|0\leq i<l,0\leq j<m\}$ can be expressed as a power of $xy$. Consequently, any polynomial in $\mathbb{F}_{2}[x,y]/(x^{l}+1,y^{m}+1)$ can be expressed in $\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$.
The rest of the proof is similar to Proposition 1 in [10]. Given the column space of $H_{X}$ is equal to
$\begin{split}\text{colsp}(H_{X})&=\{H_{X}\bm{x}|\bm{x}\in\mathbb{F}_{2}^{2lm}\}\\ &=\{A\bm{u}+B\bm{v}|\bm{u},\bm{v}\in\mathbb{F}_{2}^{lm}\},\end{split}$ (17)
it can be represented in terms of polynomials
$\begin{split}\text{colsp}(H_{X})&=\{a(\pi)u(\pi)+b(\pi)v(\pi)|\\ &u(\pi),v(\pi)\in\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)\}\end{split}$ (18)
Since $\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$ is a univariate polynomial ring, $a(\pi)\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$ and $b(\pi)\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$ are principal ideals. Thus, $\text{colsp}(H_{X})$ is an principal ideal and is generated by $g(\pi)$ and $\text{rank}(H_{X})=\dim\text{colsp}(H_{X})=lm-\deg g(\pi)$. Therefore, the dimension is given by
$\begin{split}k&=2lm-2\text{rank}(H_{X})\\ &=2lm-2(lm-\deg g(\pi))=2\deg g(\pi).\end{split}$ (19)
Data: $l,m,\tau_{d},p(\pi)$ ;
  /* $p(\pi)$ is a factor of $\pi^{lm}+1$/
Result: codes of parameters $[[2lm,k,d]]$
$C\leftarrow$ all polynomials $f(\pi)$ in $\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$ s.t. $\text{wt}(f(\pi))=3$;
$C^{\prime}\leftarrow$ all polynomials $c(\pi)$ in $C$ s.t. $c(\pi)\mod p(\pi)=0$;
$L\leftarrow$ all polynomial pairs $(a(\pi),b(\pi))$ in $C^{\prime}$ s.t. $\text{GCD}(a(\pi),b(\pi))=p(\pi)$;
$L^{\prime}\leftarrow\textnormal{{remove\_equivalent}}(L)$;
for $i\leftarrow 1$ to $|L^{\prime}|$ do
[…]
             $k\leftarrow 2lm-2\textnormal{{rank}}(H_{X})$;
             $d\leftarrow\textnormal{{distance\_upperbound}}(H_{X},H_{Z},\tau_{d})$;
       else
             continue ;
       end if
end for
Algorithm 2 An algorithm to search for BB codes with the new form of polynomials.
Coprime-BB codes can be viewed as a special case of generalized bicycle (GB) codes. In GB codes, polynomials are identified as sums of cyclic shift matrices. In contrast, coprime-BB codes identify polynomials as sums of the Kronecker products of two cyclic shift matrices with coprime dimensions. By adapting the new construction to Algorithm 1, we propose Algorithm 2. The latter algorithm results in a significantly reduced search space as only coprimes and codes with desired $k$ are being considered. Using Algorithm 2, we found a number of interesting coprime-BB codes, which are shown in Table 2. The results are not exhaustive as numerous polynomials meet the condition $\text{GCD}(a(\pi),b(\pi))=p(\pi)$. As the algorithm outputs codes with the same $k$, the results shown were obtained by running the search algorithm for a short duration and selecting the output code with the highest $d$. Additionally, we visualized the newly found codes along with the codes proposed in [4] in Fig. 1 as a 3D plot of the $n,k,d$ dimensions. “Better codes” are in the red subspace, which have higher $k$ and $d$ relative to their lower $n$ values, i.e., their cost in physical qubits in balanced by the logical qubits and sufficiently large distance for error correction.
We observe that the $[[126,12,10]]$ code offers the highest $d$ and one of the highest $k$, which is comparable to the $[[144,12,12]]$ “gross” code proposed in [4], making it suitable for scenarios where error rates are moderate to high, and qubit resources are ample. We also noted that in [10], the author found a [[$126,12,10$]] with different polynomials, i.e., we are not the first to discover this particular code, yet our search algorithm identified it based on a different basis. The $[[42,6,6]]$ code has a low $d$ but provides the highest rate in these codes and is best suited for scenarios with low error rates or where hardware resources are limited. The $[[70,6,8]]$ code offers a balanced trade-off between error correction capability and code length, making it a versatile option for environments with moderate physical error rates and resource constraints.
Figure 1: The visualization of newly found codes and codes previously found ones in [4].
Table 2: Some novel coprime-BB codes found by Algorithm 2
$l$ $m$ $a(\pi)$ $b(\pi)$ $[[n,k,d]]$
3 5 $1+\pi+\pi^{2}$ $\pi+\pi^{3}+\pi^{8}$ $[[30,4,6]]$
3 7 $1+\pi^{2}+\pi^{3}$ $\pi+\pi^{3}+\pi^{11}$ $[[42,6,6]]$
5 7 $1+\pi+\pi^{5}$ $1+\pi+\pi^{12}$ $[[70,6,8]]$
2 27 $\pi^{2}+\pi^{5}+\pi^{44}$ $\pi^{8}+\pi^{14}+\pi^{47}$ $[[108,12,6]]$
7 9 $1+\pi+\pi^{58}$ $\pi^{3}+\pi^{16}+\pi^{44}$ $[[126,12,10]]$
5 Simulation Results
In this section, we evaluate the error rates of the newly found codes through theoretical analysis and numerical simulation. We begin by illustrating the performance of the codes under the code-capacity model, followed by a Monte Carol simulation based on the circuit-based noise model. The simulation results are gathered under two stopping criteria: (1) when the number of logical errors reaches 100 or (2) when the total number of simulated samples reaches 10 000, whichever occurs later.
5.1 Code-Capacity Model
In the code-capacity model, we assume that all gate operations and measurements are perfect. Under this assumption, and using an ideal syndrome decoder to decode an $[[n,k,d]]$ code, all errors with a weight less than or equal to $t=\lfloor(d-1)/2\rfloor$ can be perfectly decoded. The logical error rate $p_{L}$ as a function of the physical error rate $p$ is
$p_{L}=1-\bigg{(}\sum_{i=0}^{t}\binom{n}{i}p^{i}(1-p)^{n-i}\bigg{)}.$ (20)
The logical error rates (y-axis) of proposed coprime-BB codes and the $[[144,12,12]]$ BB code are shown for different physical error rates (x-axis) in Fig. 2.
As expected, codes with a larger distance $d$ generally exhibit better performance than those with a smaller $d$. When comparing two codes with the same distance $d$, the shorter code tends to have a lower logical error rate. This is evident when we compare the $[[144,12,12]]$ BB code with the $[[126,12,10]]$ coprime-BB code, where the former significantly outperforms the latter.
Figure 2: Theoretic logical error rates if all errors with weights less than $\lfloor(d-1)/2\rfloor$ are corrected.
5.2 Circuit-Level Noise Model
In the circuit-level noise model, was assume that errors can happen during any operation except for the classical part. We adopt the common practice as described in [4], where the syndrome extraction is performed in multiple cycles. In each cycle, sources of noise are applied uniformly on measurement and gate operations (including idle gates) with a certain probability. The number of cycles is the same as the distance of the code, and after performing the desired cycles, the syndrome history is fed into the decoder to estimate the final error. Fig. 3 shows the logical error rate (y-axis) of selected codes for different error rates of physical qubit operations (x-axis). The logical error rate per cycle can be calculated by dividing the logical error rate by the number of cycles for the code. The decoder used in the simulation is BP-OSD [9], with a maximum of 10,000 min-sum (MS) iterations, variable scaling factor, and the “OSD_CS” method of order 7.
Similar to the code-capacity model, codes with higher $d$ achieve a lower logical error rate per cycle. However, we observe that the performance gap between the $[[144,12,12]]$ code and the $[[126,12,10]]$ code is much smaller than in Fig. 2. This reduced gap is likely due to the circuit-level distance $d_{c}$ being smaller than the code distance $d$, as errors can propagate through the circuit. Our Monte Carlo simulations indicate that the $[[144,12,12]]$ code has a circuit-level distance of $d_{c}\leq 10$, while the $[[126,12,10]]$ code has $d_{c}\leq 9$.
Figure 3: Logical error rates of selected codes under circuit-level noise model.
6 Related Work
Panteleev and Kalachev proposed a BP-OSD decoder and discovered the [[$126,12,10$]] code [10]. More recently, various constructions based on BB codes have been proposed to achieve different objectives. Voss et al. [11] expanded the concept of BB codes by introducing an additional type of indeterminate variable, leading to the creation of trivariate bicycle (TB) codes. These new codes reduce the weight of stabilizers from 6 to 4-5, making them more practical for hardware implementation. However, it is important to note that some of these codes exhibit a lower rate or distance compared to BB codes of similar length.
Eberhardt et al. [12] investigated the algebraic structure of BB codes and uncovered certain symmetry properties. These properties allow for the explicit construction of logical operators and certain fault-tolerant gates for BB codes. Notably, the paper proposes a $[[108,16,6]]$ code and a $[[162,24,6]]$ code, which are precisely two and three times the parameters $n,k$ of our $[[54,8,6]]$ BB code, respectively. The relationships among these three codes remain an area for future research.
Shaw et al. [13] proposed a “morphing” circuit design for syndrome extraction on BB codes. The proposed circuit design has only six rounds of CNOT gates instead of seven in [4]. By applying the proposed circuit, the authors discovered a new family of BB codes, including codes that have the same $[[n,k,d]]$ parameters as [4]. Besides, the authors also provide a sufficient condition for the circuit to be applied to the other two-block group algebra (2BGA) codes.
In addition, various BB codes have been proposed in papers related to the circuit realization of BB codes [5, 7]. Our work differentiates itself from these approaches by introducing a general and efficient algorithm for searching BB codes in the form described in Eq. (7), which is also the form of codes proposed by the related works. Besides that, we proposed a novel algorithm that allows us to search for BB codes with the desired dimensions, which we call coprime-BB codes. The coprime-BB codes also generalized the form Eq. (7) by allowing mixed terms, making them unattainable through searches limited to the polynomial form in Eq. (7).
7 Conclusion and Future Work
We developed algorithms for fast numerical searches for the discovery of BB codes. Furthermore, we proposed a novel construction of BB codes by choosing a factor polynomial from $\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$, where $l$ and $m$ are coprime integers. The new construction enables us to know the rate of BB codes before constructing them. Simulations have been done to illustrate the error rates of the newly found codes. The $[[126,12,10]]$ code achieves a slightly higher error rate than the $[[144,12,12]]$ code with fewer qubits. However, a challenge that remains is the practical implementation of these codes, specifically in mapping them onto quantum architectures that are constrained by the limitations of current quantum device technologies.